Logistics Solutions provides order fulfillment services for dot.com merchants. The company maintains warehouses that stock items carried by its dot.com clients. When a client receives an order from a customer, the order is forwarded to Logistics Solutions, which pulls the item from storage, packs it, and ships it to the customer. The company uses a predetermined variable overhead rate based on direct labor-hours.

In the most recent month, 150,000 items were shipped to customers using 5,900 direct labor-hours. The company incurred a total of $18,290 in variable overhead costs.

According to the company’s standards, 0.03 direct labor-hours are required to fulfill an order for one item and the variable overhead rate is $3.15 per direct labor-hour.

Required:

1. What is the standard labor-hours allowed (SH) to ship 150,000 items to customers?

2. What is the standard variable overhead cost allowed (SH × SR) to ship 150,000 items to customers?

3. What is the variable overhead spending variance?

4. What is the variable overhead rate variance and the variable overhead efficiency variance?
Answer #1

Standard hours allowed = 150,000*0.03

= 4500

Standard voh cost allowed =4500*3.15

= 14175

Voh spending variance = 14175-18290

= 4115 U

Voh rate variance = (SR-AR) *AH

= (3.15 - 18290/5900)*5900

= 295 F

Voh efficiency variance = (SH-AH) *SR

= (4500-5900)*3.15

= 4410 U